TEXT   25

wfd-log

Guest on 12th June 2022 03:32:04 AM

  1. 10183 -- First WFD run with TOHM/ERP/LIP
  2. 10240+(un po) -- Should be stable to here.  Various cable problems etc...
  3.  
  4. Things go nuts with the detector....
  5.  
  6. 10297 10330 - back to stable running
  7.  
  8. Stopped for following work:
  9.  
  10. We pointed ST tracks into boxes and found various problems.  We *fixed*
  11. LIP/PHRASE cables 3B11/3B14, 3B15/16, 6T5/6T8.  
  12. There is a swapped PHRASE cable on 4T7/4T8.  We have *not fixed* this.
  13. LIP/PHRASE tanks 1E08, 3T12, 3E04, 5B05/06 are DEAD.  SM2 no PHRASE
  14.  
  15. TOHM SM1 was unplugged to trigger but not STOP master.  Caused WFDs to
  16. be stopped between events which resulted in 20% muon loss. *fixed*
  17.  
  18. SM1 need one more USER_PROCEDURE 1 to set thresholds. added to
  19. gut16.lis1.  Also deleted the USER_PROCEDURE after threshold setting.
  20.  
  21. Replaced card #124 with #168 in 5B09-16 to fix a cared with ped
  22. always=100
  23.  
  24. 10331 - 10345
  25.  
  26. For the previous runs we noticed 3w05 was unplugged from the WFDs
  27.  
  28. 4T1-8 were accidently left unplugged on LIP.
  29.  
  30. LIP on SM1 has 20% inefficiency because there was a new JORWAY
  31. controller and when the TOHM got plugged in on SM1 it was clearing the
  32. LIP.  protective fix put in CAMAC list.
  33.  
  34. did some 1/2 hour data size runs.
  35.  
  36. 10346 - CSPAM/HIPT/FMT test run
  37. 10347 - normal test run
  38. 10348 - HIPT/FMT test run
  39.  
  40. FMT added to normal running.
  41.  
  42. 10349 -

Raw Paste


Login or Register to edit or fork this paste. It's free.